# On Oscillatory Integral Operators Satisfying the cinematic curvature condition

### v1

@@ Abstract.
We study the oscillatory integral operators satisfying the cinematic curvature condition. First, we formulate a conjecture for this class of operators, motivated by certain necessary conditions arising from counterexamples. We then establish an estimate for these operators by combining Wolff’s two-ends reduction with refined decoupling inequalities.
@@ 1.1. Statement of results
for all integrable $f\colon B^{n-1}\to\mathbb{C}$. In this case $T^{\lambda}$ is said to be an oscillatory integral operator satisfying the cinematic curvature condition (or cone condition).
@@
(1.1) $\|T^{\lambda}f\|_{L^{p}(\mathbb{R}^{n})}\lesssim_{\phi,a}\|f\|_{L^{p}(\mathbb{R}^{n-1})}$
@@
Observe that the Knapp example indicates that estimate (1.1) may fail when $n=3$ and $p\leq 4$. Consequently, the range (1.2) is sharp in the case $n=3$.
We denote by $B_{r}^{m}$ any ball of radius $r$ in $\mathbb{R}^{m}$. By an $\varepsilon$-removal trick, Theorem 1.1 can be reduced to
@@
$\|T^{\lambda}f\|_{L^{p}(B^{n}_{R})}\lesssim_{\phi,a,\varepsilon,p}R^{\varepsilon}\|f\|_{L^{p}(\mathbb{R}^{n-1})}$
@@
Although estimates for oscillatory integral operators satisfying the cinematic curvature condition may be regarded as a generalized extension of cone restriction estimates, the polynomial partitioning techniques that have recently proved effective in the study of the cone restriction problem (see [11] and [14]) do not seem to apply directly in this setting. The main obstacle lies in the fact that the transverse equidistribution estimate, which plays a crucial role in the polynomial partitioning framework, may not hold for oscillatory integral operators under cinematic curvature.
Our proof of Proposition 1.2 relies primarily on Wolff’s two-ends reduction and refined decoupling inequalities. This methodological framework has recently been applied in the study of local smoothing estimates by Gan-Wu [4] and in the the restriction problem by Wang–Wu [13]. Accordingly, our argument may be viewed as similar in spirit to their approach in the present setting of oscillatory integral operators satisfying the cinematic curvature condition.
In addition, we provide a counterexample showing that when the estimate (1.1) may fail. This example is inspired by the work of Bourgain-Guth [2] and arises from the Kakeya compression phenomenon, where tubes are arranged within a neighborhood of a low-dimensional set. A detailed discussion of this example is given in Section 2.
Theorem 1.3.
Suppose $T^{\lambda}$ is an oscillatory integral operator satisfying the cinematic curvature condition. The estimate (1.1) may fail when
(1.4) $p<2+\frac{8}{3n-1}$ if $n$ is odd;
$p<2+\frac{8}{3n-2}$ if $n$ is even.
@@ 2. Necessary Conditions
In this section, we examine an example of oscillatory integral operators that satisfy the cinematic curvature condition, with the aim of demonstrating the necessity of the $p$-range stated in Theorem 1.3. The construction is motivated by the work of Bourgain–Guth [2] (see also [10] [15] [7]) and is closely related to the Kakeya compression phenomenon, in which a large collection of tubes is efficiently packed into a neighbourhood of a low-dimensional set.
@@
Suppose the estimate (1.1)) holds. Since $\|f\|_{L^{p}(\mathbb{R}^{n-1})}\sim 1$, (2.9) gives
@@
This gives Theorem 1.3. Observe that when $n\leq 5$, the Knapp example yields a larger range of exponents for which estimate (1.1) may fail. Consequently, the necessary conditions established in this paper are not optimal. It is reasonable to anticipate that a sharper counterexample will be identified in subsequent work.
@@ 3.1. Some Basic Reductions
$\|T^{\lambda}f\|_{L^{p}(B^{n}_{R}(0))}\lesssim_{\phi,\varepsilon,p}R^{\varepsilon}\|f\|_{L^{p}(B^{n-1})}$
holds uniformly for $f\in\mathcal{S}(\mathbb{R}^{n-1})$ with ${\rm supp}f\subset\mathbb{A}$, $1\leq R\leq\lambda^{1-\varepsilon}$ and $\varepsilon>0$, provided that the amplitude $a$ satisfies (3.1) and the exponent $p$ lies in the range (1.3).
@@ 5. $(\textup{M}_{\bf A})$
Suppose $T^{\lambda}$ is an oscillatory integral operator satisfying the cinematic curvature condition of type $\mathbf{1}$. Assume (3.1), then for all $p$ satisfies (1.2), the estimate
$\|T^{\lambda}f\|_{L^{p}(B_{R}^{n}(0))}\lesssim_{\varepsilon,p,\phi}R^{\varepsilon}\|f\|_{L^{p}(\mathbb{R}^{n-1})}$
@@ 3.2. Wave packet decomposition
Fix $0<\varepsilon\ll 1$. In this subsection, we recall the wave packet decomposition for a function $f\in\mathcal{S}(\mathbb{R}^{n-1})$ at scale $R$ satisfying $1\ll R\leq\lambda^{1-\varepsilon}$, following the presentation in [4, Section 3.3].
@@
[4, Definition 3.3] For $(x_{n},\xi)\in[-1,1]\times\mathbb{A}$, let $\Phi(u,x_{n};\xi)$ be the unique solution to
@@
[4, Definition 3.4] Let $U\subset\mathbb{R}^{m}$ be a box of dimensions $a_{1}\times\dots\times a_{m}$. Define $U^{*}$ to be the dual box of $U$ centered at the origin of dimensions $a_{1}^{-1}\times\dots\times a_{m}^{-1}$, and with edges parallel to the corresponding edges of $U$.
@@
[4, Remark 3.6] For a slab $U\subset\mathbb{R}^{n-1}$ of dimensions $Rs^{2}\times Rs\times\dots\times Rs$ with $s\in[R^{-1/2},1]$ and a constant $C>0$, define $CU$, the non-isotropic $C$-dilation of $U$, to be a slab of dimensions $C^{2}Rs^{2}\times CRs\times\dots\times CRs$ with the same center as $U$.
@@
[4, Lemma 3.12] $(T^{\lambda}f_{T}){\bf{1}}_{B_{R}^{n}(0)}$ is essentially supported in $T$. In other words,
@@
[4, Definition 3.5] Given $\xi_{0}\in\mathbb{A}$ and a box $V\subset B_{R}^{n-1}(0)$, define
@@
[4, Lemma 3.9] Let $s\in[R^{-1/2},1]$ and $\xi_{0}\in\mathbb{A}$. Let $V\subset\mathbb{R}^{n-1}$ be a slab of dimensions $Rs^{2}\times Rs\times\dots\times Rs$. Then for $|t|\leq R$, each $t$-slice $\Gamma_{V}(\xi_{0};R)\cap\{x_{n}=t\}$ is comparable to a $Rs^{2}\times Rs\times\dots\times Rs$-slab.
@@ 3.3. Some simple estimates
[9, Theorem 1.1] Suppose $q\geq 2$. Let $S^{\lambda}$ be an oscillatory integral operator defined as follows:
@@ 3.4. Mixed Norm
Suppose $T^{\lambda}$ is an oscillatory integral operator satisfying the cinematic curvature condition of type $\mathbf{1}$. Assume (3.1) and (3.2),then for all $p$ satisfies (1.2), the estimate
$\|T^{\lambda}f\|_{L^{p}(B_{R}^{n}(0))}\lesssim_{\varepsilon,p,\phi}R^{\varepsilon}\|f\|_{L^{2}(\mathbb{R}^{n-1})}^{\frac{2}{p}}\max_{\theta\in\Theta_{R^{-1/2}}}\|f\|_{L^{2}_{\mathrm{avg}}(3\tilde{\theta})}^{1-\frac{2}{p}}$
@@
then we have
$\max_{\theta\in\Theta_{R^{-1/2}}}\|f\|_{L_{\mathrm{avg}}^{2}(3\tilde{\theta})}^{2}+{\rm RapDec}(R)\|f\|_{2}^{2}\gtrsim u^{2}R^{\frac{n-2}{4}}|\mathbb{T}|^{1/2}$
@@
(3.12) $|\mathbb{T}\cap\mathcal{T}_{\theta^{\prime}}|\gtrsim\max\{1,|\mathbb{T}|R^{-\frac{n-2}{2}}\}\geq\ |\mathbb{T}|^{1/2}R^{-\frac{n-2}{4}}$
@@ 3.5. Lorentz rescaling
We now state the main result of this subsection in the form of a lemma. The proof follows the framework developed in [1, Section 2.5] and [4, Section 5.1].
@@
Let $n\geq 3$. Suppose there exists a constant $C>0$ such that, for every function $\widetilde{g}$ with ${\rm supp}\ \widetilde{g}\subset\mathbb{A}$ and every oscillatory integral operator $T^{\lambda/K^{2}}$ satisfying the cinematic curvature condition of type $\mathbf{1}$, the estimate
(3.14) $\|T^{\lambda/K^{2}}\widetilde{g}\|_{L^{p}(B_{R/K^{2}}^{n})}^{p}\leq C\,\|\widetilde{g}\|_{2}^{2}\max_{\theta\in\Theta_{KR^{-1/2}}}\|\widetilde{g}\|_{L^{2}_{\mathrm{avg}}(3\tilde{\theta})}^{\,p-2}+{\rm RapDec}(R/K^{2})\,\|\widetilde{g}\|_{2}^{p}$
@@
$\|T^{\lambda}g\|_{L^{p}(B_{R}^{n}(0))}^{p}\lesssim_{\phi}C\,K^{\,2-(n-2)(p-2)}\,\|g\|_{2}^{2}\max_{\theta\in\Theta_{R^{-1/2}}}\|g\|_{L^{2}_{\mathrm{avg}}(3\tilde{\theta})}^{\,p-2}+{\rm RapDec}(R)\,\|g\|_{2}^{p}.$
@@
(3.19) $\|\widetilde{T}^{\lambda/K^{2}}\tilde{g_{\Box}}\|_{L^{p}(B^{n}_{R/K^{2},\Box})}^{p}\lesssim_{\phi}C\,\|\tilde{g_{\Box}}\|_{2}^{2}\max_{\theta\in\Theta_{KR^{-1/2}}}\|\tilde{g_{\Box}}\|_{L^{2}_{\mathrm{avg}}(3\tilde{\theta})}^{\,p-2}\\ +{\rm RapDec}(R/K^{2})\,\|\tilde{g_{\Box}}\|_{2}^{p}$
@@
$\|T^{\lambda}g\|_{L^{p}(B_{R}^{n}(0))}^{p}\lesssim_{\phi}K^{2-(n-2)(p-2)}\max_{\theta\in\Theta_{R^{-1/2}}}\|g\|_{L^{2}_{\mathrm{avg}}(3\tilde{\theta})}^{p-2}\sum_{\Box}\|g_{\Box}\|_{2}^{2}+{\rm RapDec}(R)\,\|g\|_{2}^{p}.$
Invoking Lemma 3.7, we conclude the proof. ∎
@@ 3.6. Simplification of phase function
From [4, Proof of Lemma 5.4], one can reduce Proposition 3.14 to the case
@@
Suppose $T^{\lambda}$ is an oscillatory integral operator satisfying the cinematic curvature condition of type $\mathbf{1}$. Assume (3.1), (3.2), (3.22) and (3.23), then for all $p$ satisfies (1.2), the estimate
$\|T^{\lambda}f\|_{L^{p}(B_{R}^{n}(0))}\lesssim_{\varepsilon,p,\phi}R^{\varepsilon}\|f\|_{L^{2}(\mathbb{R}^{n-1})}^{\frac{2}{p}}\max_{\theta\in\Theta_{R^{-1/2}}}\|f\|_{L^{2}_{\mathrm{avg}}(3\tilde{\theta})}^{1-\frac{2}{p}}$

### v2

@@ Abstract.
We establish an almost sharp $L^{r}\mapsto L^{p}$ estimate for oscillatory integral operators satisfying the cinematic curvature condition. The proof combines Wolff’s two-ends reduction with refined decoupling inequalities.
@@ 1.1. Statement of results
for all integrable $f\colon B^{n-1}\to\mathbb{C}$. In this case $T^{\lambda}$ is said to be an oscillatory integral operator satisfying the cinematic curvature condition.
@@
(1.1) $\|T^{\lambda}f\|_{L^{p}(\mathbb{R}^{n})}\lesssim_{\phi,a}\|f\|_{L^{r}(\mathbb{R}^{n-1})}$
@@
and
(1.3) $\frac{r}{r-1}\leq\frac{n-2}{n}p$
The range of $r$ in (1.3) is sharp, as shown by the Knapp example.
Prior to this work, the best-known result on this problem was due to Schippa [13], who verified that the estimate (1.1) holds for
$p>\begin{cases}4&\text{if $n=3$},\\[4.0pt] 2+\frac{8}{3n-3}&\text{if $n>3$ odd},\\[4.0pt] 2+\frac{8}{3n-4}&\text{if $n>3$ even}.\end{cases}$
and $r=p$ using polynomial partitioning. However, for the full range of exponents $r$ satisfying (1.3), that result only implies that the estimate (1.1) holds for
$p>2+\frac{4}{n-1}$
We denote by $B_{r}^{m}$ any ball of radius $r$ in $\mathbb{R}^{m}$. By a $\varepsilon$-removal trick together with bilinear interpolation (see [12, 13] for details), Theorem 1.1 can be reduced to
@@
Suppose $T^{\lambda}$ is an oscillatory integral operator satisfying the cinematic curvature condition. Then the estimate (1.1) may fail when
$\|T^{\lambda}f\|_{L^{p}(B^{n}_{R})}\lesssim_{\phi,a,\varepsilon}R^{\varepsilon}\|f\|_{L^{r}(\mathbb{R}^{n-1})}$
@@
and $r$ satisfies (1.3).
Our proof of Proposition 1.2 relies primarily on Wolff’s two-ends reduction and refined decoupling inequalities. This methodological framework has recently been applied in the study of local smoothing estimates by Gan-Wu [5] and in the the restriction problem by Wang–Wu [15]. Accordingly, our argument may be viewed as similar in spirit to their approach in the present setting of oscillatory integral operators satisfying the cinematic curvature condition. Compared with the polynomial partitioning approach [12, 16, 13], this framework performs substantially better for $L^{r}\mapsto L^{p}$ estimates. For instance, Gao–Wu–Xi [4] recently obtained an almost sharp $L^{r}\mapsto L^{p}$ bound for Hörmander-type operators.
In addition, we provide a counterexample showing that when the estimate (1.1) may fail. After an earlier version of this manuscript was made available, we learned that a related example had previously been given in [13]. For completeness, we include our version here. This example is inspired by the work of Bourgain-Guth [2] and arises from the Kakeya compression phenomenon, where tubes are arranged within a neighborhood of a low-dimensional set. A detailed discussion of this example is given in 2.
Proposition 1.3.
Suppose $T^{\lambda}$ is an oscillatory integral operator satisfying the cinematic curvature condition, then the estimate (1.1) may fail when
(1.5) $p<\begin{cases}2+\frac{8}{3n-1}&\text{if $n$ is odd},\\[4.0pt] 2+\frac{8}{3n-2}&\text{if $n$ is even}.\end{cases}$
and $r$ satisfies (1.3).
Comparing (1.2) with (1.3), we see that our result is almost sharp, up to an $O(n^{-2})$ error in the $p$-exponent.
@@ 2. Necessary Conditions
In this section, we examine an example of oscillatory integral operators that satisfy the cinematic curvature condition, with the aim of proving Proposition 1.3. The construction is motivated by the work of Bourgain–Guth [2] (see also [11] [17] [8]) and is closely related to the Kakeya compression phenomenon, in which a large collection of tubes is efficiently packed into a neighbourhood of a low-dimensional set.
After an earlier version of this manuscript was circulated, we became aware that a related example had previously appeared in [13]. For completeness, we include our version here with full details.
@@
Suppose $r$ satisfies (1.3) and the estimate (1.1)) holds. Since $\|f\|_{L^{r}(\mathbb{R}^{n-1})}\sim 1$, (2.9) gives
@@
This gives Proposition 1.3.
@@ 3.1. Some Basic Reductions
$\|T^{\lambda}f\|_{L^{p}(B^{n}_{R}(0))}\lesssim_{\phi,\varepsilon,p}R^{\varepsilon}\|f\|_{L^{r}(B^{n-1})}$
holds uniformly for $f\in\mathcal{S}(\mathbb{R}^{n-1})$ with ${\rm supp}f\subset\mathbb{A}$, $1\leq R\leq\lambda^{1-\varepsilon}$ and $\varepsilon>0$, provided that the amplitude $a$ satisfies (3.1) and the exponents $p,r$ satisfy (1.4) and (1.3).
@@ 5. $(\textup{M}_{\bf A})$
Suppose $T^{\lambda}$ is an oscillatory integral operator satisfying the cinematic curvature condition of type $\mathbf{1}$. Assume (3.1), then for all $p$ satisfies (1.4) and $r$ satisfies (1.3), the estimate
$\|T^{\lambda}f\|_{L^{p}(B_{R}^{n}(0))}\lesssim_{\varepsilon,p,\phi}R^{\varepsilon}\|f\|_{L^{r}(\mathbb{R}^{n-1})}$
@@ 3.2. Wave packet decomposition
Fix $0<\varepsilon\ll 1$. In this subsection, we recall the wave packet decomposition for a function $f\in\mathcal{S}(\mathbb{R}^{n-1})$ at scale $R$ satisfying $1\ll R\leq\lambda^{1-\varepsilon}$, following the presentation in [5, Section 3.3].
@@
[5, Definition 3.3] For $(x_{n},\xi)\in[-1,1]\times\mathbb{A}$, let $\Phi(u,x_{n};\xi)$ be the unique solution to
@@
[5, Definition 3.4] Let $U\subset\mathbb{R}^{m}$ be a box of dimensions $a_{1}\times\dots\times a_{m}$. Define $U^{*}$ to be the dual box of $U$ centered at the origin of dimensions $a_{1}^{-1}\times\dots\times a_{m}^{-1}$, and with edges parallel to the corresponding edges of $U$.
@@
[5, Remark 3.6] For a slab $U\subset\mathbb{R}^{n-1}$ of dimensions $Rs^{2}\times Rs\times\dots\times Rs$ with $s\in[R^{-1/2},1]$ and a constant $C>0$, define $CU$, the non-isotropic $C$-dilation of $U$, to be a slab of dimensions $C^{2}Rs^{2}\times CRs\times\dots\times CRs$ with the same center as $U$.
@@
[5, Lemma 3.12] $(T^{\lambda}f_{T}){\bf{1}}_{B_{R}^{n}(0)}$ is essentially supported in $T$. In other words,
@@
[5, Definition 3.5] Given $\xi_{0}\in\mathbb{A}$ and a box $V\subset B_{R}^{n-1}(0)$, define
@@
[5, Lemma 3.9] Let $s\in[R^{-1/2},1]$ and $\xi_{0}\in\mathbb{A}$. Let $V\subset\mathbb{R}^{n-1}$ be a slab of dimensions $Rs^{2}\times Rs\times\dots\times Rs$. Then for $|t|\leq R$, each $t$-slice $\Gamma_{V}(\xi_{0};R)\cap\{x_{n}=t\}$ is comparable to a $Rs^{2}\times Rs\times\dots\times Rs$-slab.
@@ 3.3. Some simple estimates
[10, Theorem 1.1] Suppose $q\geq 2$. Let $S^{\lambda}$ be an oscillatory integral operator defined as follows:
@@ 3.4. Mixed Norm
Suppose $T^{\lambda}$ is an oscillatory integral operator satisfying the cinematic curvature condition of type $\mathbf{1}$. Assume (3.1) and (3.2),then for all $p$ satisfies (1.4) and $r$ satisfies (1.3), the estimate
$\|T^{\lambda}f\|_{L^{p}(B_{R}^{n}(0))}\lesssim_{\varepsilon,p,\phi}R^{\varepsilon}\|f\|_{L^{2}(\mathbb{R}^{n-1})}^{\frac{2}{r}}\max_{\theta\in\Theta_{R^{-1/2}}}\|f\|_{L^{2}_{\mathrm{avg}}(3\tilde{\theta})}^{1-\frac{2}{r}}$
@@
then for all $\gamma\in[0,1]$ we have
$\max_{\theta\in\Theta_{R^{-1/2}}}\|f\|_{L_{\mathrm{avg}}^{2}(3\tilde{\theta})}^{2}+{\rm RapDec}(R)\|f\|_{2}^{2}\gtrsim u^{2}R^{\frac{n-2}{2}(1-\gamma)}|\mathbb{T}|^{\gamma}$
@@
(3.12) $|\mathbb{T}\cap\mathcal{T}_{\theta^{\prime}}|\gtrsim\max\{1,|\mathbb{T}|R^{-\frac{n-2}{2}}\}\geq\ |\mathbb{T}|^{\gamma}R^{-\gamma\frac{n-2}{2}}$
@@ 3.5. Lorentz rescaling
We now state the main result of this subsection in the form of a lemma. The proof follows the framework developed in [1, Section 2.5] and [5, Section 5.1].
@@
Let $n\geq 3$ and $2<r\leq p$. Suppose there exists a constant $C>0$ such that, for every function $\widetilde{g}$ with ${\rm supp}\ \widetilde{g}\subset\mathbb{A}$ and every oscillatory integral operator $T^{\lambda/K^{2}}$ satisfying the cinematic curvature condition of type $\mathbf{1}$, the estimate
(3.14) $\|T^{\lambda/K^{2}}\widetilde{g}\|_{L^{p}(B_{R/K^{2}}^{n})}\leq C\,\|\widetilde{g}\|_{2}^{\frac{2}{r}}\max_{\theta\in\Theta_{KR^{-1/2}}}\|\widetilde{g}\|_{L^{2}_{\mathrm{avg}}(3\tilde{\theta})}^{1-\frac{2}{r}}+{\rm RapDec}(R/K^{2})\,\|\widetilde{g}\|_{2}$
@@
$\|T^{\lambda}g\|_{L^{p}(B_{R}^{n}(0))}\lesssim_{\phi}C\,K^{\frac{n}{p}-(n-2)(1-\frac{1}{r})}\,\|g\|_{2}^{\frac{2}{r}}\max_{\theta\in\Theta_{R^{-1/2}}}\|g\|_{L^{2}_{\mathrm{avg}}(3\tilde{\theta})}^{1-\frac{2}{r}}+{\rm RapDec}(R)\,\|g\|_{2}.$
@@
(3.19) $\|\widetilde{T}^{\lambda/K^{2}}\tilde{g_{\Box}}\|_{L^{p}(B^{n}_{R/K^{2},\Box})}^{p}\lesssim_{\phi}C\,\|\tilde{g_{\Box}}\|_{2}^{\frac{2p}{r}}\max_{\theta\in\Theta_{KR^{-1/2}}}\|\tilde{g_{\Box}}\|_{L^{2}_{\mathrm{avg}}(3\tilde{\theta})}^{\,p-\frac{2p}{r}}\\ +{\rm RapDec}(R/K^{2})\,\|\tilde{g_{\Box}}\|_{2}^{p}$
@@
$\|T^{\lambda}g\|_{L^{p}(B_{R}^{n}(0))}^{p}\lesssim_{\phi}K^{n-(n-2)(p-\frac{p}{r})}\max_{\theta\in\Theta_{R^{-1/2}}}\|g\|_{L^{2}_{\mathrm{avg}}(3\tilde{\theta})}^{p-\frac{2p}{r}}\sum_{\Box}\|g_{\Box}\|_{2}^{\frac{2p}{r}}+{\rm RapDec}(R)\,\|g\|_{2}^{p}.$
Since $r\leq q$, we can now conclude the proof invoking Lemma 3.7. ∎
@@ 3.6. Simplification of phase function
From [5, Proof of Lemma 5.4], one can reduce Proposition 3.14 to the case
@@
Suppose $T^{\lambda}$ is an oscillatory integral operator satisfying the cinematic curvature condition of type $\mathbf{1}$. Assume (3.1), (3.2), (3.22) and (3.23), then for all $p$ satisfies (1.4) and $r$ satisfies (1.3), the estimate
$\|T^{\lambda}f\|_{L^{p}(B_{R}^{n}(0))}\lesssim_{\varepsilon,p,\phi}R^{\varepsilon}\|f\|_{L^{2}(\mathbb{R}^{n-1})}^{\frac{2}{r}}\max_{\theta\in\Theta_{R^{-1/2}}}\|f\|_{L^{2}_{\mathrm{avg}}(3\tilde{\theta})}^{1-\frac{2}{r}}$